Assume Gillette Corporation will pay an annual dividend of $ 0.66$0.66 one year from now. Analysts expect this dividend to grow at 12.7 %12.7% per year thereafter until the 44th year.​ Thereafter, growth will level off at 2.4 %2.4% per year. According to the​ dividend-discount model, what is the value of a share of Gillette stock if the​ firm's equity cost of capital is 7.6 %7.6%​?
Answer #1

Expected Dividend in 1 year, D1 = $0.66

Growth rate of 12.70% until fourth year and a constant growth rate (g) of 2.40% thereafter.

D2 = $0.6600 * 1.1270 = $0.7438
D3 = $0.7438 * 1.1270 = $0.8383
D4 = $0.8383 * 1.1270 = $0.9448
D5 = $0.9448 * 1.0240 = $0.9675

Cost of Capital, rs = 7.60%

P4 = D5 / (rs - g)
P4 = $0.9675 / (0.0760 - 0.0240)
P4 = $18.6058

P0 = $0.66/1.076 + $0.7438/1.076^2 + $0.8383/1.076^3 + $0.9448/1.076^4 + $18.6058/1.076^4
P0 = $16.51

So, current stock price is $16.51